OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2) (Points : 1) -11 0 11 undefined you can give me the answer for an example (because i have many other questions like this) but if you do can you also explain how you got it. you also dont have to give the answer if you explain!

OpenStudy (anonymous):

slope = (Y1 - Y0)/(X1 - X0), for given points P1=(X1, Y1) and P0=(X0, Y0)

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

she knows slope already

OpenStudy (anonymous):

As you can see (5, 9) and (5, -2) will occur a division by zero, answer: Undefined

OpenStudy (texaschic101):

there is a formula.. slope(m) = (y2 - y1) / (x2 - x1) (5,9) x1 = 5 and y1 = 9 (5,-2) x2 = 5 and y2 = -2 now we sub slope(m) = (-2 - 9) / (5 - 5) slope(m) = -11/0 when there is a 0 in the denominator, it is a vertical line and the slope is undefined. If there would have been a 0 in the numerator, it would have been a horizontal line with a slope of 0. VNOX = vertical line, no slope or undefined, x = a constant (0 in the denominator) HOY = horizontal line, 0 slope, y = a constant (0 in the numerator)